What is the GFM for N2O6? (N = 14.0g, O = 16.0g)



1
Expert's answer
2021-06-16T05:32:30-0400

GFM for N2O6 is

M(N2O6)=2M(N)+6M(O)=2*14g/mole+6*16g/mole=124 g/mole